OpenStudy (anonymous):

PLEASE HELP!! Find the coordinates of three locations or items that, when connected together, form a triangle. List the coordinates. Find the length and slope of each side of your triangle. Classify your triangle based on the side lengths and slopes found as either equilateral, right, or isosceles. If your triangle does not fit one of those categories, explain why not. These are the coordinates (13,18) , (16,15) and (20,20)

OpenStudy (anonymous):

I already know its an isosceles triangle and stuff but i need to find the length and slope of each side please help!

OpenStudy (notamathgenius):

Just graph them.

jigglypuff314 (jigglypuff314):

to find the distance between two points, you can use the distance formula \[distance formula: \sqrt{(x _{2}-x _{1})^{2} + (y _{2}-y _{1})^{2}} \]when given points (x1, y1) and (x2, y2) to find the slope between two points, you can use the slope formula \[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} } \]

OpenStudy (linda3):

@jigglypuff314 got to it before i did! :)

OpenStudy (anonymous):

Thank you!! @jigglypuff314

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (anonymous):

I have one more question! with the distance formula are you square rooting it to? @jigglypuff314

OpenStudy (anonymous):

do you square root it

jigglypuff314 (jigglypuff314):

yes :) the distance formula is based off the pythagorean theroem c^2 = a^2 + b^2 c = sqrt (a^2 + b^2) so you would sqrt

OpenStudy (anonymous):

Oh okay makes sense now. Thanks again! @jigglypuff314

jigglypuff314 (jigglypuff314):

glad I could help :)

OpenStudy (anonymous):

for the distance formula do you do x times 2? @jigglypuff314

jigglypuff314 (jigglypuff314):

no sorry for the confusion, that "2" is to tell the difference between the x value in the second point versus the x value in the first point so like given points (1,2) and (3,4) x1 = 1 y1 = 2 x2 = 3 y2 = 4

OpenStudy (anonymous):

for points (13,18) and (16,15) ..... it would be: x1= 13 y1=18 x2= 16 y2= 15 so when i set it up it should be (16-13)2 + (15-18)2 would that be then 6+ 6 = 12 square root it so the answer would be 3.4? @jigglypuff314

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

yes, that's what I got with those numbers :) 3.464 which could be 3.5 if you wanted to round up?